Circular No. 8086

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003bj, 2003bk, 2003bl, AND 2003bm Three new apparent supernovae have been found on unfiltered CCD

images — SN 2003bk and SN 2003bl by B. Swift, D. Weisz, and W. Li (LOTOSS; cf. IAUC 8075), and SN 2003bm by T. Boles (cf. IAUC 8073):

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003bk Feb. 28.5 12^{h}22^{m}42.^{s}14 + 9^{o}20′01′′.2 17.0 1′′.9 W, 4′′.2 N
2003bl Mar. 2.5 13 57 30.65 + 6 05 36.4 18.4 15′′.1 E, 12′′.6 S
2003bm Mar. 2.89 8 07 22.87 +40 23 44.3 18.0 16′′.6 E, 10′′.1 S

Additional approximate magnitudes by the respective discoverers: SN 2003bk in NGC 4316, Jan. 31.5 UT, [18.5; Mar. 2.3, 17.0. SN 2003bl in NGC 5374, 2002 June 13.3, [19.0; Mar. 3.5, 18.4. SN 2003bm in UGC 4226, Feb. 20, [19.0; Mar. 3.989, 18.0. SN 2003bm is also absent from Palomar Sky Survey red and blue plates.

M. Phillips and M. Hamuy, Carnegie Observatories, report that a spectrum (range 380–930 nm) of SN 2003bk, obtained on Mar. 3.28 UT with the Las Campanas Dupont 2.5-m telescope (+ WFCCD), shows it to be a highly reddened type-II supernova having a red continuum with a strong Hα line exhibiting a P-Cyg profile. The minimum of the Hα absorption yields an expansion velocity of 7600 km/s, assuming the NED recession velocity of 1252 km/s for the host galaxy. A spectrum of SN 2003bl, taken on Mar. 3.35, shows it to be a young type-II supernova with a blue continuum, weak P-Cyg Balmer lines, and He I at 587.6 nm; the minimum of the Hβ absorption yields an expansion velocty of 6300 km/s, assuming the NED recession velocity of 4295 km/s for the host galaxy. Hamuy adds that a spectrum of SN 2003bj (cf. IAUC 8085), obtained as above on Mar. 2.30 by M. Roth and N. Morrell, shows it to be a young type-II supernova having a blue continuum with weak Hα and Hβ P-Cyg lines and a well-defined He I 587.6-nm line.

NOVA IN M81 Further to their reports on IAUC 8085, A. V. Filippenko and R.

Chornock also write that inspection of CCD spectra, obtained (as related on IAUC 8084) on Feb. 27 UT, shows that the ‘possible nova in M81’ that was reported on IAUC 8069 is indeed a nova. Strong, double-peaked hydrogen Balmer emission lines are visible. Many Fe II, Ca II, and O I lines also appear in emission. An estimated magnitude of the object from the TV guider is R = 22–23.

2003 March 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green